OpenStudy (anonymous):

Circumference of the circle A is 24pie and circumference of the circle B is 12 pie . Which the difference between the areas of A and B?

OpenStudy (anonymous):

108pi

OpenStudy (anonymous):

\[Circumference_A=(2\pi)r_A=24\pi\]\[Circumference_B=(2\pi)r_B=12\pi\]Work out the radius in each, then use the formula: \[Area_A=\frac{1}{2}(2\pi)r_A^2\]\[Area_B=\frac{1}{2}(2\pi)r_B^2\]

OpenStudy (anonymous):

Using the circumference formulas, we know that the radius of circle A is 12 and the radius of circle B is 6. Therefore, A's area would be 144pi and the area of B would be 36pi. Then, just subtract these values to find the answer.